OpenStudy (anonymous):

Can someone show me the step to solve this please. I've looked at the examples that my teach provided and in my book but there are none for this type of problem. Thanks Express in terms of i. Root -225 + root -49

OpenStudy (anonymous):

any time you have a sum of a negative number in a square root you write i so for the firstone it is 15i + 7i= 22i

OpenStudy (anonymous):

We have \[\sqrt{-1} = i\]now \[\sqrt{-225} = \sqrt{225}\times \sqrt{-1} = 15i\]similarly, \[\sqrt{-49} = 7i\]. Add them and you have your answer.. :)

OpenStudy (anonymous):

K thank you. Know I can take it from there :)

OpenStudy (anonymous):

Both of you thanks :)